OpenStudy (anonymous):

Chemical A freezes at 13 degrees lower than Chemical B if Chemical A freezes at -36 degrees celsius at what temperature does Chemical B freezes?

OpenStudy (anonymous):

Let b = B's freezing temperature.

OpenStudy (anonymous):

Type the equation, then press Enter. (Omit the degree symbol.) ^thats how it says in my computer

OpenStudy (lalaly):

a=b-13 -36=b-13 b=-23

OpenStudy (anonymous):

13°C - 36°C = -23°C ^Lalaly thanks :)